A hockey puck (mass = 2 kg) leaves the players stick (moving to the left) with a speed of 10 m/s and slides on the ice before coming to rest. The coefficient of friction between the puck and the ice is 0.4 What is the normal force on the puck? Submit Answer Tries 0/2 What is the friction force exerted on the puck due to the ice? Submit Answer Tries 0/2 What is the magnitude of the acceleration of the puck due to this friction force? m/s Submit Answer Tries 0/2 How long will the puck slide after leaving the players stick? (what amount of time will it be sliding) Submit Answer Tries 0/2 How far will the puck slide after leaving the players stick? Submit Answer Tries 0/2
Answer #1

in vertical, F_net = FN - Fg = 0

FN = 2 x 9.8 = 19.6 N ......Ans


-----------

fk = uk N = 0.4 x 19.6 = 7.84 N ....Ans

(to the right)

---------------

a = fk/m = 3.92 m/s^2 .....Ans

(to the right)

----------------

vf = vi + a t

0 = - 10 + 3.92 t

t = 2.55 sec .....Ans

--------


vf^2 - vi^2 = 2 a d

0^2 - 10^2 = 2(-3.92) d

d = 12.8 m ........ANs